OpenStudy (anonymous):

best answer gets medal (d - 5)( ? ) = - 25 A. 5 - d B. 5 + 5d C. d - 5 D. d + 5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

im guessing c. :/

OpenStudy (skullpatrol):

So you are guessing that $$(d - 5)( d - 5 ) = - 25$$

OpenStudy (anonymous):

hhmm...

OpenStudy (skullpatrol):

The question is asking what goes where the question mark is to make the expression equal to -25?

OpenStudy (anonymous):

Cannot be determine.

OpenStudy (anonymous):

sorry i messed up its actually(d - 5)( ? ) =d^2 - 25 A.5 - d B.5 + 5d C.d - 5 D.d + 5

OpenStudy (skullpatrol):

Ok, try to fill in the question mark ( ? ) with one of the choices. ? means what goes here.

OpenStudy (skullpatrol):

$$(d - 5)( ? ) =d^2 - 25$$

OpenStudy (anonymous):

yas and im still thinking its c

OpenStudy (anonymous):

d+5

OpenStudy (skullpatrol):

So you still think $$(d - 5)( d - 5 ) =d^2 - 25$$

OpenStudy (anonymous):

c or d

OpenStudy (skullpatrol):

Do you know the FOIL method of multiplying binomials?

OpenStudy (anonymous):

it is d+5....(d+5)*(d-5)= D^2-25

OpenStudy (anonymous):

so.. D.

OpenStudy (skullpatrol):

@Sourav16 Code of Conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

ok. then how do you do the foil process ?

OpenStudy (skullpatrol):

@johnathan.... Do you know the method?

OpenStudy (anonymous):

i didn't say the answer..i jus mentioned the formula...ok fyn...sorry....:(

OpenStudy (skullpatrol):

1. Multiply the First terms of the binomials 2. Multiply the first term of each binomial by the last term of the other and add these products. 3. multiply the last terms of the binomials

OpenStudy (skullpatrol):